OpenStudy (anonymous):

43 = 16 + 3x Slove For x

OpenStudy (anonymous):

43 = 16 + 3x 43 - 16 = 3x 27 = 3x Divide 27 / 3 to get your answer!

OpenStudy (anonymous):

Did you get x = 9? That's your answer.

OpenStudy (anonymous):

You can give medals by clicking "Best Response".

OpenStudy (anonymous):

3x=43-16 3x=27 x=9